# leftmargin=*

### v3

@@ Abstract
Many humanoid and multi-legged robots are controlled in positions rather than in torques, preventing direct control of contact forces, and hampering their ability to create multiple contacts to enhance their balance, such as placing a hand on a wall or a handrail. This paper introduces the SEIKO (Sequential Equilibrium Inverse Kinematic Optimization) pipeline, drawing inspiration from flexibility models used in serial elastic actuators to indirectly control contact forces on traditional position-controlled robots. SEIKO formulates whole-body retargeting from Cartesian commands and admittance control using two quadratic programs solved in real time. We validated our pipeline with experiments on the real, full-scale humanoid robot Talos in various multi-contact scenarios, including pushing tasks, far-reaching tasks, stair climbing, and stepping on sloped surfaces. This work opens the possibility of stable, contact-rich behaviors while getting around many of the challenges of torque-controlled robots. Code and videos are available at https://hucebot.github.io/seiko_controller_website/.
@@
Whole Body Admittance Control, Multi-Contact, Teleoperation, Joint Flexibility, Humanoid Robot.
@@ I Introduction
Humans often use additional contact points to enhance their stability, for instance, using a handrail or a wall when walking, or to extend their reach, for instance, to grasp an object that is too far forward. While humanoid robots would benefit from a similar strategy, current robots minimize the number of contacts and use them only for feet and required interactions with the environment, such as pushing a button [1].
The primary challenge in controlling multi-contact lies in the redundancy of force distribution resulting from closed kinematic chains [2]. For a given posture with several contacts, there exists an infinity of ways to distribute force among them. For instance, a humanoid with both hands on a table can apply more or less force to the hands without any visible change in joint position.
To explicitly regulate forces, most work on multi-contact whole-body control relies on torque-controlled robots with inverse dynamics controllers [3, 4, 5]. Unfortunately, inverse dynamics is highly sensitive to model and calibration errors, and identifying models for humanoids is particularly challenging [6]. Perfect identification of environment’s properties is generally not possible. This is why most deployed robots use position control, which is simpler and more reliable [7], but lacks direct control authority over contact force, thus hindering the exploitation of multi-contact strategies.
In this paper, we present a control pipeline (Fig. 1) designed to regulate contact forces using position-controlled robots. Drawing inspiration from series elastic actuators [8], our main idea is to leverage the “flexibility” that stems from mechanical parts that bend with high load but also from the non-perfect joint position tracking with PID controllers, which act as “elastic elements”. While it is not possible to compute the contact force from the posture alone [2], the flexibility will make the forces converge to a unique equilibrium. We model this phenomenon to link position commands to force distribution. To invert this relation, that links force distributions to joint positions, we formulate it as a Quadratic Programming (QP) problem to leverage fast QP solvers.
@@
We conducted experiments on the Talos humanoid robot [9], equipped with powerful arms but known for significant hip mechanical flexibility [10]. Our control pipeline is compatible with commands from autonomous planners and teleoperation, with a focus on the latter in this study. Well-suited for teleoperation, our method is robustness against operator errors related to awareness and embodiment challenges. Unlike most existing methods, our approach enables motions close to feasibility boundaries (both in term of kinematic, balance and torque limits), allowing to fully exploit the capabilities of the hardware.
@@
  • An SQP formulation that computes posture deflection and joint command correction, accounting for joint flexibility in multi-contact quasi-static conditions.
@@ II Related Work
Our previous works [11, 12, 13] explored teleoperation and retargeting for feasible multi-contact tasks on simulated humanoids and hardware bimanual manipulators. While torque-controlled robots and inverse dynamic controllers were used for regulating contact wrenches, this new work introduces a whole-body admittance controller to address position-controlled robots, enhancing robustness to model errors.
Many works have studied teleoperation of complex robots with floating bases [14], but few have explicitly addressed multi-contact scenarios. [15] demonstrated multi-contact teleoperation with operator whole-body tracking on HRP-4, including retargeting and position control. [16] introduced a dedicated human-robot interface for multi-contact teleoperation on Valkyrie, employing joint impedance control. However, both approaches lack explicit regulation of contact forces.
On torque-controlled robots, multi-contact tasks have been explored both in simulation [3] and with real humanoids [4, 5]. Whole-body inverse dynamic controllers are used for direct regulation of contact wrenches and internal forces.
Joint impedance control provides a robust alternative to pure torque control, accommodating model errors while enabling force commands. [17, 18] showcased multi-contact setups on CENTAURO and COMAN+ robots, employing a QP-based whole-body inverse kinematics for postural control. A second QP calculates contact force references from quasi-static assumption, integrating these references as torque feedforward term in the low-level joint impedance scheme.
On position-controlled robots, contact force regulation is frequently ignored. [19] demonstrated balance stabilization through feedback laws applied prior to inverse kinematics, and [20] showcased ladder climbing on HRP-2 using an inverse dynamic controller optimizing joint accelerations, integrated twice to obtain position commands. While purely kinematic control is effective when the robot is far from feasibility limits, it inherently lacks control over the robot’s full configuration.
Foot force difference control from [21] influenced a common approach for regulating contact forces: applying an admittance scheme on the effector’s Cartesian position normal to the contact surface [22, 23, 24, 25]. This scheme implicitly relies on joint impedance or flexibility, often requiring ad-hoc feedback laws on foot height, ankle joints, and CoM.
The idea of explicitly modeling torques produced from position-controlled actuators, proposed by [26] and studied in [27], has been applied to multi-contact tasks on the Walk-Man robot [28, 29]. Similar to our approach, they differentiate the quasi-static equilibrium but their method uses pseudo-inverses and does not consider retargeting nor constraints.
@@
The closest related work is [30], showcasing multi-contact tasks on the position-controlled HRP-2 humanoid. They differentiate the quasi-static equilibrium and use elastic joint models, but their method solves a cascade of QP problems, while our formulation is unified. Their purely reactive control architecture, lacking feedforward terms and retargeted references, is more sensitive to noise and violations of the quasi-static assumption. In contrast, our method allows faster motions and assumes actual joint positions cannot be measured but are estimated by the model, accommodating robots with mechanical flexibility like the Talos robot.
@@ III Problem Definition
Quasi-static robot configurations are defined by postural positions, joint torques, and contact wrenches $\bm{q},\bm{\tau},\bm{\lambda}$. For position-controlled robots, control inputs only consist of joint position commands $\bm{\theta}^{\text{cmd}}$. The whole-body retargeting stage (Fig. 1, [11]) provides a stream of desired quasi-static configurations $\bm{q}^{\text{d}},\bm{\tau}^{\text{d}},\bm{\lambda}^{\text{d}}$ expected to be feasible.
@@
  • Multi-contact tasks exhibit redundancy in both kinematics and contact wrench distribution, akin to the Grasp matrix’s nullspace in manipulation [31]).
@@
  • To ensure safety, physical limits must be enforced such as balance, joint kinematics, actuator torque limits, and contact stability conditions [32] prohibiting pulling, sliding, tilting.
  • To apply the controller to hardware, it must be robust to model errors and violations of simplifying assumptions.
@@ IV-A Main Idea
According to rigid body theory in multi-contact [2, 31], the contact wrenches of an ideal infinitely stiff mechanical system are non-unique and lie in a redundant nullspace. Real systems, however, always exhibit inherent flexibility: the structure slightly bends, and both the deflected posture and contact wrenches uniquely evolve towards the configuration minimizing overall elastic energy. Therefore, given constant joint position commands, the mapping taking into account flexibility $\bm{\theta}^{\text{cmd}}\mapsto(\bm{q}^{\text{flex}},\bm{\lambda}^{\text{flex}})$ is unique and well-defined. Our approach models and predicts this whole-body non-linear deflection effect, utilizing it for the control of contact wrenches.
Specifically, we differentiate and linearize the deflection effect to consider how contact wrenches change with variations in joint position commands through the Jacobian matrix $\frac{\partial\bm{\lambda}^{\text{flex}}}{\partial\bm{\theta}^{\text{cmd}}}(\bm{q}^{\text{flex}},\bm{\lambda}^{\text{flex}},\bm{\theta}^{\text{cmd}})$. Instead of directly inverting this Jacobian matrix, we formulate the control problem as a Quadratic Programming (QP) which solves for position command changes and optimizes multiple objectives, similar to task space inverse dynamic approaches. We explicitly model the system’s flexibility by treating each robot joint as a spring, encompassing both internal actuator impedance and mechanical flexibilities.
@@ IV-B Overall Architecture
The set of effectors that may come into contact with the environment is pre-defined. Each effector’s state is either: “enabled”, standing for fixed and in contact transmitting forces and torques to the environment, or “disabled”, indicating that it is free to move and is commanded by the operator. Our formulation handles both plane contacts (6 DoFs, e.g., feet) and point contacts (3 DoFs, e.g., hands).
An external planner or human operator provides commands as input to the Retargeting stage: (i) Cartesian pose $\bm{X}^{\text{op}}$ or velocity $\bm{\nu}^{\text{op}}$ commands for each free (disabled) effector, (ii) a Boolean signal that manually triggers the transition between contact states, and (iii) an optional “pushing mode” enabling explicit control of the normal force of a specific enabled contact. Our method does not plan contact sequencing, relying on external decisions for contact stances and sequence.
@@ IV-C Equilibrium Equation and Flexibility Model
which is non-linear in $\bm{q}$. The equilibrium equation is linearized by considering small variations of the configuration $\left(\bm{q}+\Delta\bm{q},\bm{\lambda}+\Delta\bm{\lambda},\bm{\tau}+\Delta\bm{\tau}\right)$. The differentiation is written:
@@
while neglecting second order terms.
@@
Note that link flexibility can also be modeled in a similar manner by introducing passive joints without actuation. Its differentiated expression is written:
@@
The differentiated equilibrium equation (2) combined with flexibility model (4) is linear w.r.t. configuration changes:
@@
$\begin{bmatrix}\frac{\partial\bm{g}}{\partial\bm{q}}(\bm{q}^{\text{flex}})-\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q}^{\text{flex}})\bm{\lambda}^{\text{flex}}\right)-\bm{S}\bm{K}\bm{S^{\prime}}&|\leavevmode\nobreak\ -\bm{J}(\bm{q}^{\text{flex}})^{\mathsf{T}}\end{bmatrix},$
@@
where $\bm{T}_{\text{B}},\bm{t}_{\text{B}}$ stands for the floating base rows and $\bm{T}_{\text{J}},\bm{t}_{\text{J}}$ for the joint rows.
@@ IV-D SEIKO Retargeting
This section summarizes the SEIKO Retargeting method developed in [11, 12].
@@
$\bm{X}^{\text{target}}_{\text{off}}=\mathsf{filtering}\left(\bm{X}_{\text{ref}}(t)\oplus\bm{X}^{\text{op}}\right)$ (8)
@@
where $\bm{X}_{\text{ref}}\in SE(3)$ is a reference pose that integrates velocity commands at each time step. It allows the Cartesian pose command to be expressed relative to this reference. The filtering process incorporates a smoothing low-pass filter and enforces signal’s velocity and acceleration limits through time-optimal bang-bang trajectory planning. We also constrain $\bm{X}_{\text{ref}}$ within a radius of $\bm{X}^{\text{d}}_{\text{off}}$ to prevents the reference pose to windup when the retargeted motion is saturated by the feasibility constraints.
@@
differentiated equilibrium equation (2) (9g)
@@
$C_{\text{contact}}(\bm{\lambda}^{\text{d}})\Delta\bm{\lambda}^{\text{d}}+c_{\text{contact}}(\bm{\lambda}^{\text{d}})\geq\bm{0}$ (9k)
@@
Equality constraints enforce the equilibrium equation (9g) and ensure enabled contacts are fixed (9h). Inequality constraints include joint position limits (9i), joint torque limits (9j), and contact stability conditions (9k) considering unilaterality, friction pyramid, and center of pressure (see [32]). Additional constraints involve limits on joint changes (9l) and contact wrench changes (9m).
We enhanced the contact switching procedure compared to prior work. To remove a contact, we instantly increase the weight of the wrench penalty task to a very high value and use joint velocity $\dot{\bm{\theta}}^{\text{max}}$ and wrench velocity $\dot{\bm{\lambda}}^{\text{max}}$ limits to ensure a smooth transition. When the integrated desired wrench falls below a small threshold, the contact is removed. Enabling a contact is straightforward, as it doesn’t require any special considerations, thanks to these limits.

### v4

# Multi-Contact Whole-Body Force Control for Position-Controlled Robots

@@ Abstract
Many humanoid and multi-legged robots are controlled in positions rather than in torques, which prevents direct control of contact forces, and hampers their ability to create multiple contacts to enhance their balance, such as placing a hand on a wall or a handrail. This paper introduces the SEIKO (Sequential Equilibrium Inverse Kinematic Optimization) pipeline, and proposes a unified formulation that exploits an explicit model of flexibility to indirectly control contact forces on traditional position-controlled robots. SEIKO formulates whole-body retargeting from Cartesian commands and admittance control using two quadratic programs solved in real time. Our pipeline is validated with experiments on the real, full-scale humanoid robot Talos in various multi-contact scenarios, including pushing tasks, far-reaching tasks, stair climbing, and stepping on sloped surfaces. Code and videos are available at: https://hucebot.github.io/seiko_controller_website/
@@
Multi-Contact Whole-Body Motion Planning and Control, Whole-Body Motion Planning and Control, Humanoid Robot Systems.
@@ I Introduction
Humans often use additional contact points to enhance their stability, for instance, by using a handrail or a wall when walking, or to extend their reach, for instance, when grasping a distant object. While humanoid robots would benefit from a similar strategy, current robots minimize the number of contacts and use them only for feet and required interactions with the environment, such as pushing a button [1].
The primary challenge in controlling multi-contact lies in the redundancy of force distribution resulting from closed kinematic chains [2]. For a given posture with several contacts, there are infinite ways to distribute force among them. For instance, a humanoid with both hands on a table can apply more or less force to the hands without any visible change in joint position.
To regulate forces, most prior studies on multi-contact whole-body control rely on torque-controlled robots with inverse dynamics controllers [3, 4, 5]. Unfortunately, inverse dynamics is highly sensitive to model and calibration errors, and identifying models for humanoids is particularly challenging [6]. Perfect identification of environment’s properties is generally not possible. This is why most deployed robots use position control, which is simpler and more reliable [7], but it lacks direct control authority over contact forces.
Prior work on position-controlled robots [8, 9, 10, 11] has often regulated contact forces indirectly using various forms of admittance schemes applied independently to each effector. While effective in many scenarios, this strategy may lack robustness in challenging situations near physical limits or with significant model errors. This is due to its heuristic nature, which lacks theoretical grounding and fails to consider the whole-body effect of postural changes on contact forces.
Our main idea is to exploit the robot’s non-rigidity to explicitly model the relationship between joint position commands and contact forces. Flexibility arises from either non-observable mechanical structural bending or internal impedance of non-ideal joint position control. We present a control pipeline (Fig. 1) designed to regulate contact forces on position-controlled robots. Our approach offers a novel unified whole-body formulation using optimization-based Quadratic Programming (QP) to leverage fast QP solvers.
@@
We conducted experiments on the Talos humanoid robot [12], equipped with powerful arms but known for significant hip mechanical flexibility [13]. Our control pipeline is compatible with commands from autonomous planners and teleoperation, with a focus on the latter in this study. Well-suited for teleoperation, our method is robust against operator errors related to awareness and embodiment challenges. Unlike most existing methods, our approach enables motions close to feasibility boundaries (both in term of kinematic, balance, and torque limits), allowing full exploitation of the capabilities of the hardware.
@@
  • A Sequential QP (SQP) formulation that computes posture deflection and joint command correction, accounting for joint flexibility in multi-contact quasi-static conditions.
@@ II Related Work
Multi-contact tasks have been studied in-depth on humanoid robots with torque control [3, 4, 5], where contact forces are directly regulated using whole-body inverse dynamic controllers. But torque control relies on an accurate model of the robot’s dynamic, which is challenging to identify [6, 13] and lacks robustness. Joint impedance control [14, 15] offers a more robust alternative to torque control, but still requires modeling the actuators to be able to specify torque feedforward references.
Although [16] demonstrated ladder climbing on a position-controlled robot without regulating contact forces, it is essential to regulate these forces. Doing so enables pushing tasks, smooth contact transitions, and enhances robustness for motions near system limits, where stability margins are reduced. Most studies using position-controlled robots [9, 10, 11, 17] regulate contact forces indirectly through methods named “effector admittance”, “foot force difference control”, or “damping control” all based on the same principle introduced in [8]: an admittance feedback law is applied to each effector to adjust its Cartesian pose reference. For example, to reduce the force measured on the hand, these approaches will retract its desired position away from the contact surface. However, because it is expected that the hand remains in contact, these approaches implicitly rely on flexibility without explicitly considering it. In contrast, SEIKO Controller models the whole-body flexibility, providing a more grounded formulation. This also enables the accounting of postural changes on the contact forces, which the effector admittance scheme cannot do.
@@
[18] proposed the idea of modeling torques produced by position-controlled actuators, which was further studied in [19] and applied to multi-contact in [20, 21]. Similar to our approach, they differentiate the quasi-static equilibrium but their method uses pseudo-inverses which fails at considering constraints. Furthermore, [21] also uses elastic joint models, but their method solves a cascade of several QP problems. Their purely reactive control architecture lacks feedforward terms and retargeted references, making it more sensitive to noise and violations of the quasi-static assumption. In contrast, our method is unified, allows faster motions, and does not require actual joint positions to be measured, which accommodates robots with mechanical flexibility like the Talos robot.
This work builds upon our prior work [22, 23, 24] which devised a retargeting framework for multi-contact tasks on simulated humanoids and hardware bimanual manipulators with an highlight on enforcing feasibility. Both this work and [22, 23, 24] also target teleoperation applications. While many studies have investigated the teleoperation of complex robots with floating bases [25], fewer have explicitly addressed multi-contact scenarios [26, 27]. In contrast to these, our work addresses the regulation of contact forces and demonstrates both contact switch and pushing tasks.
@@ III Problem Definition
Quasi-static robot configurations are defined by postural positions, joint torques, and contact wrenches $\bm{q},\bm{\tau},\bm{\lambda}$. For position-controlled robots, control inputs only consist of joint position commands $\bm{\theta}^{\text{cmd}}$. The whole-body retargeting stage (illustrated in Fig. 1 and proposed in previous work [22]) provides a stream of desired quasi-static configurations $\bm{q}^{\text{d}},\bm{\tau}^{\text{d}},\bm{\lambda}^{\text{d}}$ expected to be feasible.
@@
  • Multi-contact tasks exhibit redundancy in both kinematics and contact wrench distribution, akin to the Grasp matrix’s nullspace in manipulation [28].
@@
  • To ensure safety, physical limits must be enforced such as balance, joint kinematics, actuator torque limits, and contact stability conditions prohibiting pulling, sliding, tilting.
  • For application to hardware, the controller must be robust to model errors and violations of simplifying assumptions.
@@ IV-A Main Idea
According to rigid body theory in multi-contact [2, 28], the contact wrenches of an ideal infinitely stiff mechanical system are non-unique and lie in a redundant nullspace. Real systems, however, always exhibit inherent flexibility: the structure slightly bends, and both the deflected posture and contact wrenches uniquely evolve towards the configuration minimizing overall elastic energy. Therefore, given constant joint position commands, the mapping that takes into account flexibility $\bm{\theta}^{\text{cmd}}\mapsto(\bm{q}^{\text{flex}},\bm{\lambda}^{\text{flex}})$ is unique and well-defined. Our approach models and predicts this whole-body non-linear deflection effect, utilizing it for the control of contact wrenches.
Specifically, we linearize and compute derivatives of the deflection effect to consider how contact wrenches change with variations in joint position commands through the Jacobian matrix $\frac{\partial\bm{\lambda}^{\text{flex}}}{\partial\bm{\theta}^{\text{cmd}}}(\bm{q}^{\text{flex}},\bm{\lambda}^{\text{flex}},\bm{\theta}^{\text{cmd}})$. Instead of directly inverting this Jacobian matrix, we formulate the control problem as a Quadratic Programming (QP) which solves for position command changes and optimizes multiple objectives, similar to task space inverse dynamic approaches. We explicitly model the system’s flexibility by treating each robot joint as a spring, encompassing both internal actuator impedance and mechanical flexibilities.
@@ IV-B Overall Architecture
The set of effectors that may come into contact with the environment is pre-defined. Each effector’s state is either: “enabled”, standing for fixed and in contact transmitting forces and torques to the environment, or “disabled”, indicating that it is free to move and is commanded by the operator. Our formulation handles both plane contacts (6 DoFs, e.g., feet) and point contacts (3 DoFs, e.g., hands with ball shape). The full details of contact formulation are available in the supplementary material of [22]. Other types of contacts can also be easily implemented, such as full grasp contact for hand grippers or even line contact on the edge of feet.
An external planner or human operator provides commands as input to the Retargeting stage: (i) Cartesian pose $\bm{X}^{\text{op}}$ or velocity $\bm{\nu}^{\text{op}}$ commands for each free (disabled) effector, (ii) a Boolean signal that manually triggers the transition between contact states, and (iii) an optional “pushing mode” enabling explicit control of the normal force of a specific enabled contact. Our method does not plan contact sequencing, and relies on external decisions for contact stances and sequence.
@@ IV-C Equilibrium Equation and Flexibility Model
which is non-linear in $\bm{q}$. We approximate the linearization of the equilibrium equation by considering small variations of the configuration $\left(\bm{q}+\Delta\bm{q},\bm{\lambda}+\Delta\bm{\lambda},\bm{\tau}+\Delta\bm{\tau}\right)$ and partial derivatives:
@@
while neglecting second order terms (see Section -B in supplementary material for details).
@@
Note that link flexibility can also be modeled in a similar manner by introducing passive joints without actuation. Its derivative is written:
@@
The derivative-based linear approximation of the equilibrium equation (2) combined with flexibility model (4) is linear w.r.t. configuration changes:
@@
$\begin{bmatrix}\frac{\partial\bm{g}}{\partial\bm{q}}(\bm{q}^{\text{flex}})-\left(\frac{\partial\bm{J}}{\partial\bm{q}}^{\mathsf{T}}(\bm{q}^{\text{flex}})\bm{\lambda}^{\text{flex}}\right)+\bm{S}\bm{K}\bm{S^{\prime}}&|\leavevmode\nobreak\ -\bm{J}(\bm{q}^{\text{flex}})^{\mathsf{T}}\end{bmatrix},$
@@
where $\bm{T}_{\text{B}},\bm{t}_{\text{B}}$ refer to the first 6 rows representing the floating base and $\bm{T}_{\text{J}},\bm{t}_{\text{J}}$ the remaining $n$ joint rows.
@@ IV-D SEIKO Retargeting
This section summarizes the SEIKO Retargeting method developed in [22, 23]. From this previous work, Section -C in supplementary material provides further explanation on how balance is enforced.
@@
$\bm{X}^{\text{target}}_{\text{off}}=\mathsf{filtering}\left(\bm{X}_{\text{ref}}(t)\bm{X}^{\text{op}}\right),\leavevmode\nobreak\ \bm{X}_{\text{ref}}(0)=\bm{X}^{\text{read}},$ (8)
@@
where $\bm{X}_{\text{ref}}\in SE(3)$ is a reference pose that integrates velocity commands at each time step (see [23]). $\bm{X}_{\text{ref}}$ allows the Cartesian pose command $\bm{X}^{\text{op}}$ to be expressed relative to this reference, and not in arbitrary world frame. The filtering process incorporates a smoothing low-pass filter and bounds signal’s velocity and acceleration through time-optimal bang-bang trajectory replanning [23]. With the clamping $\mathsf{boundDistance}()$, we also constrain separately the position and orientation of $\bm{X}_{\text{ref}}$ within a radius centered on $\bm{X}^{\text{d}}_{\text{off}}$ to prevent the reference pose from windup when the retargeted motion is saturated by the feasibility constraints.
@@
linearized equilibrium equation (2) (9g)
@@
$\bm{C}_{\text{contact}}\Delta\bm{\lambda}^{\text{d}}+\bm{C}_{\text{contact}}\bm{\lambda}^{\text{d}}+\bm{c}_{\text{contact}}\geq\bm{0}$ (9k)
@@
Equality constraints enforce the linearized equilibrium equation (9g) and ensure enabled contacts are fixed (9h). Inequality constraints include joint position limits (9i), joint torque limits (9j), and contact stability conditions (9k) considering unilaterality, friction pyramid, and center of pressure (see Section -D in supplementary material). Additional constraints involve limits on joint changes (9l) and contact wrench changes (9m).
Compared to our prior SEIKO Retargeting work [22], we enhanced the contact switching procedure with fewer arbitrary choices and clearer physical semantics. Details can be found in Section -E in supplementary material.